# An Asymptotically Optimal Coordinate Descent Algorithm for Learning Bayesian Networks from Gaussian Models

### v1

@@ 4.1. Convergence and Optimality Guarantees
Our analysis for optimality guarantees requires an assumption on the population model. For the set $E\subseteq\{(i,j):i,j=1,2\dots,m\}$, consider the optimization problem
@@
(5) $\Gamma^{\star}_{E}=\operatorname*{arg\,min}_{\Gamma\in\mathbb{R}^{m\times m}}\sum_{i=1}^{m}-2\log(\Gamma_{ii})+\mathrm{tr}\left(\Gamma\Gamma^{\mathrm{\scriptscriptstyle T}}{\Sigma}^{\star}\right)\quad\text{ s.t. }\quad\mathrm{support}(\Gamma)\subseteq{E}.$
Assumption 2.
@@
There exists constants $\bar{\kappa},\underline{\kappa}>0$ such that $\sigma_{\text{min}}(\Gamma^{\star}_{E})\geq\underline{\kappa}$ and $\sigma_{\text{max}}(\Gamma^{\star}_{E})\leq\bar{\kappa}$ for every $E$ where the graph $(V,E)$ is a DAG, where $\sigma_{\text{min}}(\cdot)$ and $\sigma_{\text{min}}(\cdot)$ are the smallest and largest eigenvalues respectively.
We further define $d_{\text{max}}:=\max_{i}|\{j:(j,i)\in E_{\mathrm{super}}\}|$.
Theorem 7.
@@
Let $\hat{\Gamma},\hat{\Gamma}^{\mathrm{opt}}$ be the solution of Algorithm 1 and an optimal solution of (3), respectively. Suppose Assumption 2 holds and let the regularization parameter be chosen so that $\lambda^{2}=\mathcal{O}(\log{m}/n)$ where $m$ and $n$ denote the number of nodes and number of samples, respectively. Then,
  1. (1)
@@
  2. (2)
    if $n/\log(n)\geq\mathcal{O}(m^{2}\log{m})$, with probability greater than $1-1/\mathcal{O}(n)$, we have that: $0\leq f(\hat{\Gamma})-f(\hat{\Gamma}^{\mathrm{opt}})\leq\mathcal{O}(\sqrt{d_{\text{max}}^{2}m^{4}\log{m}/n})$.
In other words, the objective value of the coordinate descent solution converges in probability to the optimal objective value as $n\to\infty$. Further, assuming the sample size $n$ is sufficiently large, with high probability, the difference in objective value is bounded by $\mathcal{O}(\sqrt{d_{\text{max}}^{2}m^{4}\log{m}/n})$.
Our proof relies on the following lemmas. Throughout, we let $\hat{E}$ be the support of $\hat{\Gamma}$, i.e., $\hat{E}=\{(i,j),\hat{\Gamma}_{ij}\neq 0\}$.
Lemma 8.
Let $\hat{\Gamma},\hat{\Gamma}^{\mathrm{opt}}$ be the solution of Algorithm 1 and optimal solution of (3), respectively. Then, i) for any $u,v=1,2,\dots,m,A_{uv}+2\Gamma_{uv}\hat{\Sigma}_{uu}=2(\hat{\Sigma}\Gamma)_{uv}$ where $A_{uv}$ is defined in Proposition 1. ii) if $\hat{\Gamma}_{uv}\neq 0$, then $(\hat{\Sigma}\hat{\Gamma})_{uv}=0$, and iii) the matrix $\hat{\Gamma}\hat{\Gamma}^{\mathrm{\scriptscriptstyle T}}\hat{\Sigma}$ has ones on the diagonal.
Proof.
For $u,v=1,\ldots,m$, by the definition of $A_{uv}$, $A_{uv}+2\Gamma_{uv}\hat{\Sigma}_{uu}=2(\hat{\Sigma}\Gamma)_{uv}$, proving item i. Since any solution from Algorithm 1, $\hat{\Gamma}$ satisfies Proposition 1, for any $(u,v)\in\hat{E}$, $(4\hat{\Sigma}_{uu}\hat{\Gamma}_{uu}+A_{uu})^{2}=A_{uu}^{2}+16\hat{\Sigma}_{uu}$ and $A_{uv}=-2\hat{\Gamma}_{uv}\hat{\Sigma}_{uu}$. Combining the previous relations, we conclude that $(\hat{\Sigma}\hat{\Gamma})_{uv}=0$. Therefore, for any $(u,v)\in\hat{E}$, we have $\hat{\Gamma}_{uv}\not=0$ and $(\hat{\Sigma}\hat{\Gamma})_{uv}=0$, resulting in $\hat{\Gamma}_{uv}(\hat{\Sigma}\hat{\Gamma})_{uv}=0$. This proves item ii. Plugging $A_{uu}$ into the previous relations, we arrive at $\hat{\Gamma}_{uu}(\hat{\Sigma}\hat{\Gamma})_{uu}=1$. Thus, $(\hat{\Gamma}\hat{\Gamma}^{\mathrm{\scriptscriptstyle T}}\hat{\Sigma})_{ii}=\sum_{j=1}^{m}\hat{\Gamma}_{ij}(\hat{\Gamma}^{\mathrm{\scriptscriptstyle T}}\hat{\Sigma})_{ji}=\hat{\Gamma}_{ii}(\hat{\Gamma}^{\mathrm{\scriptscriptstyle T}}\hat{\Sigma})_{ii}=1$, proving item iii. ∎
Lemma 9.
@@
(6) $\hat{\Gamma}_{E}=\operatorname*{arg\,min}_{\Gamma\in\mathbb{R}^{m\times m}}\sum_{i=1}^{m}-2\log(\Gamma_{ii})+\mathrm{tr}\left(\Gamma\Gamma^{\mathrm{\scriptscriptstyle T}}\hat{\Sigma}\right)\quad\text{s.t.}\quad\mathrm{support}(\Gamma)\subseteq E.$
@@
Suppose that $4m\bar{\kappa}\|\hat{\Sigma}-\Sigma^{\star}\|_{2}\leq\min\{8\bar{\kappa}^{3}/m\bar{\kappa}^{2},{1}/{2m\underline{\kappa}}\}$ and that $\hat{\Sigma}$ is positive definite. Then, $\|\hat{\Gamma}_{E}-\Gamma^{\star}_{E}\|_{F}\leq 4m\bar{\kappa}\|\hat{\Sigma}-\Sigma^{\star}\|_{2}$.
Proof.
@@
The proof follows from standard convex analysis and Brouwer’s fixed point theorem; we provide the details below. Since $\Gamma$ follows a DAG structure, the objective of (6) can be written as: $-2\log\det(\Gamma)+\|\Gamma\hat{\Sigma}^{1/2}\|_{F}^{2}$. The KKT conditions state that there exists $Q$ with $\text{support}(Q)\cap E=\emptyset$ such that the optimal solution $\hat{\Gamma}_{E}$ of (6) satisfies $-2\hat{\Gamma}_{E}^{-1}+Q+2\hat{\Gamma}_{E}\hat{\Sigma}=0$ and $\mathrm{support}(\hat{\Gamma}_{E})\subseteq{E}$. Let $\Delta=\hat{\Gamma}_{E}-\Gamma^{\star}_{E}$. By Taylor series expansion, $\hat{\Gamma}_{E}^{-1}=(\Gamma^{\star}_{E}+\Delta)^{-1}={\Gamma^{\star}_{E}}^{-1}+{\Gamma^{\star}_{E}}^{-T}\Delta{\Gamma^{\star}_{E}}^{-1}+\mathcal{R}(\Delta),$ where $\mathcal{R}(\Delta)=2{\Gamma^{\star}_{E}}^{-1}\sum_{k=2}^{\infty}(-\Delta{\Gamma^{\star}_{E}})^{k}$. For any matrix $M\in\mathbb{R}^{m\times{m}}$, define the operator $\mathbb{I}^{\star}$ with $\mathbb{I}^{\star}(M):=2{\Gamma^{\star}_{E}}^{-T}M{\Gamma^{\star}_{E}}^{-1}+2M\Sigma^{\star}$. Let $\mathcal{K}$ be the subspace $\mathcal{K}=\{M\in\mathbb{R}^{m\times m}:\text{support}(M)\subseteq E\}$ and let $P_{\mathcal{K}}$ be the projection operator onto subspace $\mathcal{K}$ that zeros out entries of the input matrix outside of the support set $E$. From the optimality condition of (5), we have $\mathcal{P}_{\mathcal{K}}[2{\Gamma_{E}^{\star}}^{-1}-2\Gamma^{\star}_{E}\Sigma^{\star}]=0$. Then, the optimality condition of (6) can be rewritten as:
@@
Since $\hat{\Gamma}_{E}\in\mathcal{K}$ and ${\Gamma}^{\star}_{E}\in\mathcal{K}$, we have that $\Delta\in\mathcal{K}$. We use Brouwer’s theorem to obtain a bound on $\|\Delta\|_{F}$. We define an operator $J$ as $\mathcal{K}\to\mathcal{K}$:
$J(\delta)=\delta-(\mathcal{P}_{\mathcal{K}}\mathbb{I}^{\star}\mathcal{P}_{\mathcal{K}})^{-1}\left(\mathcal{P}_{\mathcal{K}}\left[\mathbb{I}^{\star}\mathcal{P}_{\mathcal{K}}(\delta)+\mathcal{R}(\delta)+H_{n}+2\delta(\hat{\Sigma}-\Sigma^{\star})\right]\right).$
Here, the operator $\mathcal{P}_{\mathcal{K}}\mathbb{I}^{\star}\mathcal{P}_{\mathcal{K}}$ is invertible since $\sigma_{\text{min}}(\mathbb{I}^{\star})=\sigma_{\text{min}}({\Gamma^{\star}_{E}}^{-1})^{2}\geq\frac{1}{\underline{\kappa}^{2}}$. Notice that any fixed point $\delta$ of $J$ satisfies the optimality condition (7). Furthermore, since the objective of (6) is strictly convex, we have that the fixed point must be unique. In other words, the unique fixed point of $J$ is given by $\Delta$. Now consider the following compact set: $\mathcal{B}_{r}=\{\delta\in\mathbb{R}^{m\times m}:\text{support}(\delta)\subseteq E,\|\delta\|_{F}\leq r\}$ for $r=4m\bar{\kappa}\|\hat{\Sigma}-\Sigma^{\star}\|_{2}$. By the assumption, $r\leq\min\{8\bar{\kappa}^{3}/m\underline{\kappa}^{2},\frac{1}{2\bar{\kappa}}\}$. Then, for every $\delta\in\mathcal{B}_{r}$, we have that: $\|\delta\Gamma^{\star}_{S}\|_{F}\leq m\bar{\kappa}r\leq 1/2$ and additionally, $\|\mathcal{R}(\delta)\|_{F}\leq 2m\|\Gamma^{\star}_{E}\|_{2}^{2}/\sigma_{\text{min}}(\Gamma^{\star}_{E})\|\delta\|_{2}^{2}\frac{1}{1-\|\delta\Gamma^{\star}_{E}\|_{2}}\leq 2m\bar{\kappa}^{2}_{2}/\underline{\kappa}r^{2}\frac{1}{1-r\bar{\kappa}}\leq 4m\bar{\kappa}^{2}_{2}/\underline{\kappa}r^{2}$. Since $\|H_{n}\|_{F}\leq 2m\|\Gamma^{\star}_{E}\|_{2}\|\hat{\Sigma}-\Sigma^{\star}\|_{2}$ and $\|G(\delta)\|_{F}\leq\frac{1}{\underline{\kappa}^{2}}[\|H_{n}\|_{F}+\|\mathcal{R}(\delta)\|_{F}+2\|\delta(\hat{\Sigma}-\Sigma^{\star})\|_{F}]$ we conclude that $\|J(\delta)\|_{F}\leq\frac{4m\bar{\kappa}^{2}r^{2}}{\underline{\kappa}^{3}}+\frac{4m\max\{\bar{\kappa},1\}}{\underline{\kappa}^{2}}\|\hat{\Sigma}-\Sigma^{\star}\|_{2}\leq r$. In other words, we have shown that $J$ maps $\mathcal{B}_{r}$ onto itself. Appealing to Brouwer’s fixed point theorem, we conclude that the fixed point must also lie inside $\mathcal{B}_{r}$. Thus, we conclude that $\|\Delta\|_{F}\leq r$. ∎
Lemma 10.
With probability greater than $1-1/\mathcal{O}(n)$, we have that: $\|\hat{\Sigma}-\Sigma^{\star}\|_{2}\leq\mathcal{O}(\sqrt{m\log(n)/n})$, $\|\hat{\Sigma}\|_{\infty}\leq 2\bar{\kappa}^{2}$, $\sigma_{\text{min}}(\hat{\Sigma})\geq\underline{\kappa}^{2}/2$, $\|\hat{\Gamma}\|_{\infty}\leq 2\bar{\kappa}$ and $\sigma_{\text{min}}(\hat{\Gamma})\geq\underline{\kappa}/2$.
Proof.
From standard Gaussian concentration results that when $n/\log(n)\geq\mathcal{O}(m)$, with probability greater than $1-\mathcal{O}(1/n)$, we have that $\|\hat{\Sigma}-\Sigma^{\star}\|_{2}\leq\mathcal{O}\allowbreak(\sqrt{m\log(n)/n})$. By Assumption 2, with probability greater than $1-\mathcal{O}(1/n)$, $\hat{\Sigma}$ is positive definite, with $\|\hat{\Sigma}\|_{\infty}\leq 2\bar{\kappa}^{2}$ and $\sigma_{\text{min}}(\hat{\Sigma})\geq\underline{\kappa}^{2}-\mathcal{O}(\sqrt{m\log(n)/n})\geq\underline{\kappa}^{2}/2$. Furthermore, appealing to Lemma 9 and that $n/\log(n)\geq\mathcal{O}(m^{3})$, $\|\hat{\Gamma}-{\Gamma}^{\star}_{\hat{E}}\|_{F}\leq\mathcal{O}(\sqrt{m^{3}\log(n)/n})$. Thus $\|\hat{\Gamma}\|_{\infty}\leq\|\Gamma^{\star}_{\hat{E}}\|_{2}+\bar{\kappa}\leq 2\bar{\kappa}$ and $\sigma_{\text{min}}(\hat{\Gamma})\geq\bar{\kappa}-\mathcal{O}(\sqrt{m\log(n)/n})\geq\underline{\kappa}/2$. ∎
Proof of Theorem 7.
Part 1). First,
@@
where the second inequality follows from $f(\hat{\Gamma}^{\mathrm{opt}})\geq\min_{\Theta}\{-\log\det(\Theta)+\mathrm{tr}(\Theta\hat{\Sigma})\}=\log\det(\hat{\Sigma})\allowbreak+m$; the equality follows from appealing to item i. of Lemma 8 to conclude that $f(\hat{\Gamma})=-\log\det(\hat{\Gamma}\hat{\Gamma}^{\mathrm{\scriptscriptstyle T}})+m+\lambda^{2}\|\hat{\Gamma}-\text{diag}(\hat{\Gamma})\|_{0}$.
@@
(8) $(\hat{\Gamma}\hat{\Gamma}^{\mathrm{\scriptscriptstyle T}}\hat{\Sigma})_{uv}=\sum_{i=1}^{m}\hat{\Gamma}_{ui}(\hat{\Sigma}\hat{\Gamma})_{vi}=\hat{\Gamma}_{uu}(\hat{\Sigma}\hat{\Gamma})_{vu}+\hat{\Gamma}_{uv}(\hat{\Sigma}\hat{\Gamma})_{vv}+\sum_{i\in F_{uv}}\hat{\Gamma}_{ui}(\hat{\Sigma}\hat{\Gamma})_{vi},$

### v3

@@ 4.2 Optimality guarantees
Let $d_{\text{max}}:=\max_{i}|\{j:(j,i)\in E_{\mathrm{super}}\}|$, the maximum degree of a node in $E_{\mathrm{super}}$. Our analysis for optimality guarantees requires an assumption on the population model. For the set $E\subseteq\{(i,j):i,j=1,2\dots,m\}$, consider the optimization problem
@@
$\Gamma^{\star}_{E}=\operatorname*{arg\,min}_{\Gamma\in\mathbb{R}^{m\times m}}\sum_{i=1}^{m}-2\log(\Gamma_{ii})+\mathrm{tr}\left(\Gamma\Gamma^{\mathrm{\scriptscriptstyle T}}{\Sigma}^{\star}\right)\quad\text{ s.t. }\quad\mathrm{support}(\Gamma)\subseteq{E}.$ (6)
Assumption 2
@@
There exists constants $\bar{\kappa},\underline{\kappa}>0$ such that $\sigma_{\text{min}}(\Gamma^{\star}_{E})\geq\underline{\kappa}$ and $\sigma_{\text{max}}(\Gamma^{\star}_{E})\leq\bar{\kappa}$ for every $E\subseteq E_{\mathrm{super}}$ where the graph $(V,E)$ is a DAG, where $\sigma_{\text{min}}(\cdot)$ and $\sigma_{\text{min}}(\cdot)$ are the smallest and largest eigenvalues respectively.
Theorem 7
@@
Let $\hat{\Gamma}$ be the solution of Algorithm 1 with any input topological ordering $O$ or initial point $\Gamma^{\mathrm{init}}$, and let $\hat{\Gamma}^{\mathrm{opt}}$ be an optimal solution of (3). Suppose Assumption 2 holds and let the regularization parameter be chosen so that $\lambda^{2}=\mathcal{O}(\log{m}/n)$ where $m$ and $n$ denote the number of nodes and number of samples, respectively. Then,
  1. 1.
@@
  2. 2.
    if $n/\log(n)\geq\mathcal{O}(m^{2}\log{m})$, with probability greater than $1-1/\mathcal{O}(n)$, we have $0\leq f(\hat{\Gamma})-f(\hat{\Gamma}^{\mathrm{opt}})\leq\mathcal{O}(\sqrt{d_{\text{max}}^{2}m^{4}\log{m}/n})$.
In other words, regardless of the input ordering $O$, the objective value of the coordinate descent solution converges in probability to the optimal objective value as $n\to\infty$. Further, assuming the sample size $n$ is sufficiently large, with high probability, the difference in objective value is bounded by $\mathcal{O}(\sqrt{d_{\text{max}}^{2}m^{4}\log{m}/n})$.
We prove Theorem 7 by showing the results for any coordinate-wise minimum of problem (3) (see Definition 5). Our proof relies on the following lemmas. Throughout, we let $\hat{E}$ be the support of $\hat{\Gamma}$, that is, $\hat{E}=\{(i,j),\hat{\Gamma}_{ij}\neq 0\}$.
Lemma 8
Let $\hat{\Gamma},\hat{\Gamma}^{\mathrm{opt}}$ be the solution of Algorithm 1 and optimal solution of (3), respectively. Then, i) for any $u,v=1,2,\dots,m,A_{uv}+2\hat{\Gamma}_{uv}\hat{\Sigma}_{uu}=2(\hat{\Sigma}\Gamma)_{uv}$ where $A_{uv}$ is defined in Proposition 3. ii) if $\hat{\Gamma}_{uv}\neq 0$, then $(\hat{\Sigma}\hat{\Gamma})_{uv}=0$, and iii) the matrix $\hat{\Gamma}\hat{\Gamma}^{\mathrm{\scriptscriptstyle T}}\hat{\Sigma}$ has ones on the diagonal. Moreover, these properties also hold for the optimal solution $\hat{\Gamma}^{\mathrm{opt}}$.
Proof [Proof of Lemma 8] For $u,v=1,\ldots,m$, by the definition of $A_{uv}$, $A_{uv}+2\Gamma_{uv}\hat{\Sigma}_{uu}=2(\hat{\Sigma}\Gamma)_{uv}$, proving item i. Since any solution from Algorithm 1, $\hat{\Gamma}$ satisfies Proposition 3, for any $(u,v)\in\hat{E}$, $(4\hat{\Sigma}_{uu}\hat{\Gamma}_{uu}+A_{uu})^{2}=A_{uu}^{2}+16\hat{\Sigma}_{uu}$ and $A_{uv}=-2\hat{\Gamma}_{uv}\hat{\Sigma}_{uu}$. Combining the previous relations, we conclude that $(\hat{\Sigma}\hat{\Gamma})_{uv}=0$. Therefore, for any $(u,v)\in\hat{E}$, we have $\hat{\Gamma}_{uv}\not=0$ and $(\hat{\Sigma}\hat{\Gamma})_{uv}=0$, resulting in $\hat{\Gamma}_{uv}(\hat{\Sigma}\hat{\Gamma})_{uv}=0$. This proves item ii. Plugging $A_{uu}$ into the previous relations, we arrive at $\hat{\Gamma}_{uu}(\hat{\Sigma}\hat{\Gamma})_{uu}=1$. Thus, $(\hat{\Gamma}\hat{\Gamma}^{\mathrm{\scriptscriptstyle T}}\hat{\Sigma})_{ii}=\sum_{j=1}^{m}\hat{\Gamma}_{ij}(\hat{\Gamma}^{\mathrm{\scriptscriptstyle T}}\hat{\Sigma})_{ji}=\hat{\Gamma}_{ii}(\hat{\Gamma}^{\mathrm{\scriptscriptstyle T}}\hat{\Sigma})_{ii}=1$, proving item iii. Since $\hat{\Gamma}^{\mathrm{opt}}$ is an optimal solution, it must satisfy the conditions stated in Proposition 3; otherwise, we could update $\hat{\Gamma}^{\mathrm{opt}}$ to achieve a better objective value. Therefore, the same arguments apply to $\hat{\Gamma}^{\mathrm{opt}}$.
Lemma 9
@@
$\hat{\Gamma}_{E}=\operatorname*{arg\,min}_{\Gamma\in\mathbb{R}^{m\times m}}\sum_{i=1}^{m}-2\log(\Gamma_{ii})+\mathrm{tr}\left(\Gamma\Gamma^{\mathrm{\scriptscriptstyle T}}\hat{\Sigma}\right)\quad\text{s.t.}\quad\mathrm{support}(\Gamma)\subseteq E.$ (7)
@@
Suppose that $4\sqrt{m}\bar{\kappa}\|\hat{\Sigma}-\Sigma^{\star}\|_{2}\leq\min\{\underline{\kappa}/(8\sqrt{m}\bar{\kappa}^{2}),\frac{1}{2\bar{\kappa}}\}$ and that $\hat{\Sigma}$ is positive definite. Then, $\|\hat{\Gamma}_{E}-\Gamma^{\star}_{E}\|_{F}\leq 4\sqrt{m}\bar{\kappa}\|\hat{\Sigma}-\Sigma^{\star}\|_{2}$.
@@
Proof [Proof of Lemma 9] The proof follows from standard convex analysis and Brouwer’s fixed point theorem; we provide the details below. Since $\Gamma$ follows a DAG structure, the objective of (7) can be written as: $2\log\det(\Gamma)+\|\Gamma\hat{\Sigma}^{1/2}\|_{F}^{2}$. The KKT conditions state that there exists $Q$ with $\text{support}(Q)\cap E=\emptyset$ such that the optimal solution $\hat{\Gamma}_{E}$ of (7) satisfies $2\hat{\Gamma}_{E}^{-1}+Q+2\hat{\Gamma}_{E}\hat{\Sigma}=0$ and $\mathrm{support}(\hat{\Gamma}_{E})\subseteq{E}$. Let $\Delta=\hat{\Gamma}_{E}-\Gamma^{\star}_{E}$. By Taylor series expansion, $\hat{\Gamma}_{E}^{-1}=(\Gamma^{\star}_{E}+\Delta)^{-1}={\Gamma^{\star}_{E}}^{-1}-{\Gamma^{\star}_{E}}^{-T}\Delta{\Gamma^{\star}_{E}}^{-1}+\mathcal{R}(\Delta),$ where $\mathcal{R}(\Delta)=2{\Gamma^{\star}_{E}}^{-1}\sum_{k=2}^{\infty}(-\Delta{\Gamma^{\star}_{E}})^{k}$. For any matrix $M\in\mathbb{R}^{m\times{m}}$, define the operator $\mathbb{I}^{\star}$ with $\mathbb{I}^{\star}(M):=2{\Gamma^{\star}_{E}}^{-T}M{\Gamma^{\star}_{E}}^{-1}+2M\Sigma^{\star}$. Let $\mathcal{K}$ be the subspace $\mathcal{K}=\{M\in\mathbb{R}^{m\times m}:\text{support}(M)\subseteq E\}$ and let $P_{\mathcal{K}}$ be the projection operator onto subspace $\mathcal{K}$ that zeros out entries of the input matrix outside of the support set $E$. From the optimality condition of (6), we have $\mathcal{P}_{\mathcal{K}}[2{\Gamma_{E}^{\star}}^{-1}-2\Gamma^{\star}_{E}\Sigma^{\star}]=0$. Then, the optimality condition of (7) can be rewritten as:
$\begin{aligned} &\mathcal{P}_{\mathcal{K}}\left[\mathbb{I}^{\star}(\Delta)+2\Delta(\hat{\Sigma}-\Sigma^{\star})-2\mathcal{R}(\Delta)+H_{n}\right]=0,\end{aligned}$ (8)
where $H_{n}=2\Gamma_{E}^{*}(\hat{\Sigma}-\Sigma^{*})$. Since $\hat{\Gamma}_{E}\in\mathcal{K}$ and ${\Gamma}^{\star}_{E}\in\mathcal{K}$, we have that $\Delta\in\mathcal{K}$. We use Brouwer’s theorem to obtain a bound on $\|\Delta\|_{F}$. We define an operator $J$ as $\mathcal{K}\to\mathcal{K}$:
$J(\delta)=\delta-(\mathcal{P}_{\mathcal{K}}\mathbb{I}^{\star}\mathcal{P}_{\mathcal{K}})^{-1}\left(\mathcal{P}_{\mathcal{K}}\left[\mathbb{I}^{\star}\mathcal{P}_{\mathcal{K}}(\delta)-2\mathcal{R}(\delta)+H_{n}+2\delta(\hat{\Sigma}-\Sigma^{\star})\right]\right).$
Here, the operator $\mathcal{P}_{\mathcal{K}}\mathbb{I}^{\star}\mathcal{P}_{\mathcal{K}}$ is invertible since $\sigma_{\text{min}}(\mathbb{I}^{\star})=\sigma_{\text{min}}({\Gamma^{\star}_{E}}^{-1})^{2}\geq\frac{1}{\bar{\kappa}^{2}}$. Notice that any fixed point $\delta$ of $J$ satisfies the optimality condition (8). Furthermore, since the objective of (7) is strictly convex, we have that the fixed point must be unique. In other words, the unique fixed point of $J$ is given by $\Delta$. Now consider the following compact set: $\mathcal{B}_{r}=\{\delta\in\mathbb{R}^{m\times m}:\text{support}(\delta)\subseteq E,\|\delta\|_{F}\leq r\}$ for $r=4\sqrt{m}\bar{\kappa}\|\hat{\Sigma}-\Sigma^{\star}\|_{2}$. By the assumption, $r\leq\min\{\underline{\kappa}/(8\sqrt{m}\bar{\kappa}^{2}),\frac{1}{2\bar{\kappa}}\}$. Then, for every $\delta\in\mathcal{B}_{r}$, we have $\|\delta\Gamma^{\star}_{E}\|_{F}\leq\bar{\kappa}r\leq 1/2$, and $\|\mathcal{R}(\delta)\|_{F}\leq 2\sqrt{m}\|\Gamma^{\star}_{E}\|_{2}^{2}/\sigma_{\text{min}}(\Gamma^{\star}_{E})\|\delta\|_{2}^{2}\frac{1}{1-\|\delta\Gamma^{\star}_{E}\|_{2}}\leq 2\sqrt{m}\bar{\kappa}^{2}r^{2}/\underline{\kappa}\frac{1}{1-r\bar{\kappa}}\leq 4\sqrt{m}\bar{\kappa}^{2}r^{2}/\underline{\kappa}$. Since $\|H_{n}\|_{F}\leq 2\sqrt{m}\|\Gamma^{\star}_{E}\|_{2}\|\hat{\Sigma}-\Sigma^{\star}\|_{2}$ and $\|J(\delta)\|_{F}\leq\bar{\kappa}^{2}[\|H_{n}\|_{F}+2\|\mathcal{R}(\delta)\|_{F}+2\|\delta(\hat{\Sigma}-\Sigma^{\star})\|_{F}]$, we conclude that $\|J(\delta)\|_{F}\leq\frac{8\sqrt{m}\bar{\kappa}^{2}r^{2}}{\underline{\kappa}}+\frac{\bar{\kappa}^{2}r}{2}+\frac{r^{2}}{2\sqrt{m}\bar{\kappa}}\leq r$. In other words, we have shown that $J$ maps $\mathcal{B}_{r}$ onto itself. Appealing to Brouwer’s fixed point theorem, we conclude that the fixed point must also lie inside $\mathcal{B}_{r}$. Thus, we conclude that $\|\Delta\|_{F}\leq r$.
Lemma 10
With probability greater than $1-1/\mathcal{O}(n)$, we have $\|\hat{\Sigma}-\Sigma^{\star}\|_{2}\leq\mathcal{O}(\sqrt{m\log(n)/n})$, $\|\hat{\Sigma}\|_{\infty}\leq 2\bar{\kappa}^{2}$, $\sigma_{\text{min}}(\hat{\Sigma})\geq\underline{\kappa}^{2}/2$, $\|\hat{\Gamma}\|_{\infty}\leq 2\bar{\kappa}$ and $\sigma_{\text{min}}(\hat{\Gamma})\geq\underline{\kappa}/2$.
Proof [Proof of Lemma 10] From standard Gaussian concentration results, when $n/\log(n)\geq\mathcal{O}(m)$, with probability greater than $1-\mathcal{O}(1/n)$, we have that $\|\hat{\Sigma}-\Sigma^{\star}\|_{2}\leq\mathcal{O}\allowbreak(\sqrt{m\log(n)/n})$. By Assumption 2, with probability greater than $1-\mathcal{O}(1/n)$, $\hat{\Sigma}$ is positive definite, with $\|\hat{\Sigma}\|_{\infty}\leq 2\bar{\kappa}^{2}$ and $\sigma_{\text{min}}(\hat{\Sigma})\geq\underline{\kappa}^{2}-\mathcal{O}(\sqrt{m\log(n)/n})\geq\underline{\kappa}^{2}/2$. Furthermore, appealing to Lemma 9, $\|\hat{\Gamma}-{\Gamma}^{\star}_{\hat{E}}\|_{F}\leq\mathcal{O}(\sqrt{m^{2}\log(n)/n})$. Thus, $\|\hat{\Gamma}\|_{\infty}\leq\|\Gamma^{\star}_{\hat{E}}\|_{2}+\bar{\kappa}\leq 2\bar{\kappa}$ and $\sigma_{\text{min}}(\hat{\Gamma})\geq\underline{\kappa}-\mathcal{O}(\sqrt{m\log(n)/n})\geq\underline{\kappa}/2$.
Proof [Proof of Theorem 7] Part 1). First, observe that by definition, $\hat{\Gamma}^{\mathrm{opt}}$ is the optimal solution of the problem
$\min\limits_{\Gamma\in\mathbb{R}^{m\times m}}f(\Gamma)\quad\text{s.t.}\quad\mathcal{G}\left(\Gamma-\mathrm{diag}\left(\Gamma\right)\right)\text{ is a DAG},$
@@
$f(\hat{\Gamma}^{\mathrm{opt}})\geq\min_{\Theta}\{-\log\det(\Theta)+\mathrm{tr}(\Theta\hat{\Sigma})\}=\log\det(\hat{\Sigma})+m.$ (9)
Now,
@@
where the second inequality follows from (9), and the equality follows from appealing to item iii. of Lemma 8 to conclude that $f(\hat{\Gamma})=-\log\det(\hat{\Gamma}\hat{\Gamma}^{\mathrm{\scriptscriptstyle T}})+m+\lambda^{2}\|\hat{\Gamma}-\text{diag}(\hat{\Gamma})\|_{0}$, because $\mathrm{tr}(\Theta\hat{\Sigma})\}=m$.
@@
$(\hat{\Gamma}\hat{\Gamma}^{\mathrm{\scriptscriptstyle T}}\hat{\Sigma})_{uv}=\sum_{i=1}^{m}\hat{\Gamma}_{ui}(\hat{\Sigma}\hat{\Gamma})_{vi}=\hat{\Gamma}_{uu}(\hat{\Sigma}\hat{\Gamma})_{vu}+\hat{\Gamma}_{uv}(\hat{\Sigma}\hat{\Gamma})_{vv}+\sum_{i\in F_{uv}}\hat{\Gamma}_{ui}(\hat{\Sigma}\hat{\Gamma})_{vi},$ (10)